OpenStudy (anonymous):

Please help. Is the following equation true? Use complete sentences to explain the properties used in making your decision. 4 log 3 x + 1/2 log 3 y – 2 log 3 z = log 3 x^4 sqrt(y)/z^2 .

OpenStudy (uri):

I will help... :D

OpenStudy (uri):

Okay, so to clarify; is it? \[ 4log_{3}x + \frac{1}{2}log_{y}- 2log_{3}x\]

OpenStudy (uri):

\[4log_{3}x + \frac{1}{2} log_{3}y -2log_{3}x\]

OpenStudy (uri):

\[\large 4log_{3}x + \frac{1}{2}log_{3}y - 2log_{3}x\]

terenzreignz (terenzreignz):

Strike three... you're out :D Anyway... \[\large 4\log_3x + \frac12\log_3y-2\log_3z\]

OpenStudy (uri):

Yaha sai niklo,ya mai kar rahi hu.

terenzreignz (terenzreignz):

^that's insensitive... and may not even belong here. Translate that or delete it~ it may or may not belong here.

OpenStudy (anonymous):

Use this property :